OpenStudy (anonymous):

Heather writes the equations below to represent two lines drawn on the coordinate plane. –6x + 18y = 0 4x – 12y = 20 After applying the linear combination method, Heather arrived at the equation 0 = 60. What conclusion can be drawn about the system of equations?

OpenStudy (anonymous):

The equation has no solution; therefore, the system of equations has no solution. The equation has a solution at (0, 60); therefore, the system of equations has a solution at (0, 60). The equation has infinite solutions; therefore, the system of equation as infinite solutions. The equation has a solution at (0, 0); therefore, the system of equations has a solution at (0, 0).

OpenStudy (anonymous):

Is it A

OpenStudy (anonymous):

@Babynini

OpenStudy (babynini):

hrm, i'm not sure. Try plugging (0, 60) into both equations and see if it comes out true.

OpenStudy (anonymous):

nope

OpenStudy (babynini):

Correct. Now try plugging in (0,0) to see if that's true.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

the first one is right but the second part is wrong

OpenStudy (babynini):

Right on, So we know it's not B or D.

OpenStudy (anonymous):

then its a

OpenStudy (babynini):

mhm! another solution to this would be that the girl in the problem reached 0 = 60 which is clearly not true. So there's no solution!

OpenStudy (anonymous):

thank u very much i have one more question

OpenStudy (anonymous):

Martin has a combination of 33 quarters and dimes worth a total of $6. Which system of linear equations can be used to find the number of quarters, q, and the number of dimes, d, Martin has? q + d = 6 25q + 10d = 33 q + d = 6 0.25q + 0.1d = 33 q + d = 33 25q + 10d = 6 q + d = 33 0.25q + 0.1d = 6

OpenStudy (babynini):

Er, which do you think? I'm not sure.

OpenStudy (babynini):

It will be on that ends in =6, I'm pretty sure.

OpenStudy (anonymous):

i think it either a or b

OpenStudy (babynini):

Sorry I can't help you out! I'm not sure at all :P Maybe try asking this question on a new question thing.

OpenStudy (anonymous):

its ok thanks for trying though :)

OpenStudy (babynini):

Sure thing ^-^

OpenStudy (babynini):

good luck!